Sally invested $3000 in two parts. One part was invested at 8% interest and the second part at 10% interest.
Her total earnings were $256 dollars.
How much was invested at 8% and how much was invested at 10%

Answer by bucky(2189)

We don't know how much money Sally invested at 8%. So let's call that amount X.
.
The remainder of the $3000 she had available to invest at 10%. So, the amount she had left from the $3000 to invest at 10% was $3000 - X.
.
The amount X was invested at 8% so the return that it earned is 8% times X, or equivalently 0.08 * X.
.
The amount left to invest at 10% was $3000 - X so the return that it earned is 10% times 3000 - X, or 0.10*(3000 - X). Doing the distributed multiplication, it earned 300 -0.10*X
.
When you add together the two earnings it is equal to $256. In equation form this is:
.
0.08X + 300 - 0.10X = 256
.
First, get rid of the 300 on the left side by subtracting 300 from both sides of the equation:
.
0.08X + 300 - 300 - 0.10X = 256 - 300
.
When you do this subtraction you are left with:
.
0.08X - 0.10X = -44
.
Combine the two terms on the left side and you have:
.
-0.02X = -44
.
Solve for X by dividing both sides by the coefficient (multiplier) of X which is -0.02:
.
X = -44/-0.02 = 2200
.
So now we know that Sally invested $2200 at 8%. The rest of her money is $3000 minus the $2200 which is $800, and it was invested at 10%
.
Check:
.
$2200 times 8% = $2200 * 0.08 = $176
.
and
.
$800 times 10% = $800 * 0.01 = $80
.
So the total earnings is $176 + $80 = $256
.
This is exactly the amount that the problem said it should be and, therefore, the answer that we got is correct. $2200 was invested and 8% and $800 was invested at 10%.
.
Hope that this helps you to understand the problem and the procedure to follow in solving it.